# BSTcontrol

### v2

@@ 3.4.3 Agentic AI-RAN
Agentic AI represents a significant, potentially revolutionary evolution for RAN management, shifting from traditional automation towards autonomous systems capable of pursuing complex goals with minimal human intervention [173, 106]. Operating under a principle of bidirectional interaction where AI empowers the network and the network enhances AI, agentic AI offers a paradigm for intelligent operation, administration, and maintenance in the complex environment of 6G networks. This approach enables systems to perceive, reason, decide, and act autonomously [176]. Unlike conventional AI reliant on predefined rules, agentic AI RAN aims to optimize objectives like resource allocation, service assurance, and energy efficiency through dynamic adaptation to the network environment, thereby more effectively supporting emerging services such as autonomous intelligent IoT and embodied intelligence applications.
@@
The core capabilities underpinning agentic AI RAN include sophisticated multi source perception leveraging channel information, sensing data, and visual inputs to understand the network state [175, 106]. Based on this perception, agents employ advanced reasoning processes, perhaps using chain of thought techniques to plan sequences of actions. They then interact with the network fabric by utilizing network function APIs as tools to execute the planned configurations. Crucially, effective decision making and planning are deeply reliant on integrated retrieval mechanisms. These provide agents dynamic access to essential knowledge bases containing 3GPP standards, historical logs, network policies, and other relevant information, enabling context aware reasoning and ensuring compliance [182].
@@
Building upon these capabilities, specific frameworks and architectural concepts are proposed to realize agentic AI-RAN. For example, AgentNet [106] envisions specialized networking ecosystems supporting heterogeneous AI agents like foundation model agents and embodied agents, facilitating efficient information exchange and coordination. Generative foundation models themselves can serve as interactive knowledge resources or be specialized into agents capable of understanding tasks, orchestrating workflows, and managing network components through well defined interfaces [175]. This agent driven architecture offers inherent simplifications, notably through control and user plane separation, and facilitates dynamic service orchestration and intent based networking, allowing high level goals to be translated into automated network adjustments [176].
Despite the great potential, deploying agentic AI in the RAN faces significant operational constraints. Energy efficiency, robust security, and stringent real time performance remain critical challenges. Continued research is thus essential to develop computationally efficient AI models, secure agent interactions, and highly adaptive AI driven processes tailored to the demanding operational realities of 6G [176].
@@ 3.4.4 Multi-Agent Collaboration
The distributed and complex nature of 6G systems necessitates the use of multi agent systems where multiple autonomous AI agents interact and coordinate to achieve shared or individual goals [183, 173]. Agentic AI networking inherently involves collaboration as agents deployed across different network locations such as user equipment edge servers or base stations, therefore, they must work together for tasks like resource allocation interference management or service orchestration [106]. Wireless distributed networks provide the necessary infrastructure platform for these multi agent interactions [183].
@@
Multi-agent reinforcement learning (MARL) is a key enabling paradigm allowing agents to learn collaborative strategies through interaction. MARL frameworks address how agents learn optimal policies considering the actions and states of others moving from centralized training and execution towards more decentralized approaches like centralized training with decentralized execution or fully decentralized training and execution which better suit the distributed nature of 6G [183]. Effective collaboration hinges on mechanisms for communication and information sharing considering network constraints. Techniques like graph enhanced MARL information bottleneck enhanced MARL and mirror learning are being explored to improve communication efficiency robustness and scalability in multi agent settings. Generative information retrieval can also support collaboration by providing agents access to shared knowledge bases or contextual information relevant to joint tasks [182].
@@
Structured workflows and specific agent roles are key to effective collaboration in practical agentic systems. For instance, LLM agents can automate physical layer tasks through a coordinated approach, with agents dedicated to task awareness, environmental observation, system configuration, and API invocation, all responding to high-level requests and environmental perception [175]. Likewise, generative simulation frameworks employ agents for simulation generation, test design, execution, and result interpretation, collaborating via a central orchestrator to iteratively refine network models [107]. Architectures like AgentNet further enhance inter-agent collaboration, supporting learning and knowledge transfer across diverse agent types [106]. Crucially, agentic AI also facilitates human-AI collaboration, augmenting human capabilities in knowledge-intensive tasks and complex operations [173].
@@
While multi-agent collaboration holds immense promise, significant challenges must be addressed to fully realize its potential, particularly in areas like 6G networks. These challenges include managing communication overhead, ensuring effective coordination despite partial observability, handling non-stationarity arising from evolving agent policies, and maintaining scalability as agent numbers grow [183]. Overcoming these hurdles and developing robust, efficient collaboration mechanisms is paramount to achieving the vision of truly intelligent and autonomous 6G networks.
@@ 4 Wireless Communications for Large AI Models
With the increasing complexity and size of AI models, deploying and training them across distributed wireless environments has become a critical challenge. This section explores key strategies and innovations in wireless communications that enable the efficient deployment, training, and operation of large AI models, as illustrated in Fig. 11. The related works are summarized in Table 4. Techniques like FL, split learning (SL), and federated split learning (FSL) offer advantages in privacy and efficiency but face challenges such as communication overhead, synchronization, and expert selection management. Furthermore, enabling massive access, which involves deploying numerous models and supporting simultaneous connections from a vast number of devices, necessitates scalable solutions often leveraging edge and distributed computing. Over-the-air computation (AirComp) can boost edge AI performance but presents signal interference and privacy issues. Ensuring robust security is also paramount, leading to the exploration of PLS, which leverages the characteristics of wireless channel to offer lightweight yet strong security mechanisms complementary to traditional cryptography, particularly crucial for the demands of 6G. Collectively, these approaches leverage wireless network capabilities to tackle scalability, optimize model performance, enhance security, and meet the growing demands of 6G networks and beyond.
Figure 11:   The outline of Section IV.
@@
Edge Intelligence [18] Deploying LLMs at the 6G mobile edge for key applications like healthcare and robotics control in cloud-based deployment. Identify key challenges and discuss enabling techniques for efficient LLMs edge training and inference.
[184] Wireless scheduling for age of information minimization at the network edge with unreliable orthogonal channels. Propose an online matching-while-learning algorithm and discuss its implementation for wireless scheduling.,
[185] Massive access management in 5G/B5G IoT networks with diverse QoS requirements, including URLLC. Propose a distributed cooperative multi-agent DRL approach for massive access.
DCML [186] Collaborative and privacy-preserving training of LLMs on decentralized private data using FL. Propose OpenFedLLM, a concise, integrated, and research-friendly framework/codebase for FL on LLMs.
[187] Fine-tuning LLMs using on geo-distributed private data at the network edge, considering device heterogeneity and dynamic channels. Propose an energy-efficient SL framework for LLM fine-tuning, minimizing training delay and server energy consumption.
[188] Fine-tuning LLMs over wireless communication networks, considering computation and communication delays. Propose the FedsLLM framework combining LoRA and federated split learning for LLM fine-tuning.
AirComp [189] Federated edge learning in B5G/6G networks to enhance communication efficiency and privacy. Provide an overview of Air-FEEL, outlining its basic principles, benefits, and challenges.
[190] Integrated sensing and edge AI for multi-agent environment perception over broadband wireless channels. Propose a framework that exploits feature sparsity and channel diversity for efficient over-the-air feature fusion.
[191] Distributed sensing where multiple edge devices using AirComp for server-based inference. Propose a framework using a configurable parameter to realize Max/Average pooling over-the-air.
PLS [128] Securing B5G/6G wireless networks in novel use cases leveraging PLS mechanisms. Provide a comprehensive review of PLS for 6G, connecting it with cryptographic concepts.
[192] Achieving native Layer-1 security in 6G communications, leveraging PLS combined with AI/ML within an O-RAN architecture. Propose 6G PLS frameworks integrating DNN decoding with shared keys and space-time coding.
@@ 4.1 Edge Intelligence
6G networks are expected to support in-network, distributed AI capabilities at the edge, facilitating collaborative machine learning across devices with varying computational resources [18, 193, 184]. However, one of the main challenges is that many edge devices have limited computing power and storage capacity. To address this, wireless technologies are evolving to support the distributed large AI models [9].
4.1.1 Collaborative Edge Computing
Collaborative edge computing enables distributed training of AI models across multiple edge devices, minimizing the need for large data transfers and reducing latency. By sharing model updates rather than raw data, this approach helps maintain privacy while optimizing communication resources [18]. Wireless networks provide the infrastructure to support this model, allowing edge devices to efficiently share updates with minimal communication overhead. Frameworks like NetGPT [100] optimize the distribution of large models between edge devices and cloud servers, ensuring that computations are allocated based on the available resources at the edge and in the cloud . This synergy enhances the training efficiency of large AI models while maintaining a balance between local edge computing and cloud-based processing.
@@ 4.1.2 Resource Allocation in Heterogeneous Systems
Effective resource allocation in wireless networks is essential for supporting large AI model training and deployment across edge devices [184]. In systems with heterogeneous devices and data, efficient allocation of computing, communication, and storage resources ensures fair distribution and optimal system performance. In wireless environments, managing communication resources is key to minimizing latency and optimizing bandwidth, especially for AI models requiring large amounts of data exchange. These approaches ensure that wireless networks can effectively support large-scale AI deployments, overcoming the limitations of edge devices and enabling real-time, distributed AI model training.
4.1.3 Massive Access for Large AI Model
Massive access for large AI model deployment refers to deploying numerous AI models within a network and supporting a large number of devices accessing these models simultaneously. This process involves multiple technologies and methods, including edge computing, distributed computing, and large-scale parallel processing, to ensure that the system can efficiently handle a large number of requests and data transmissions.
For massive access of devices in IoT, researchers propose a joint energy-efficient subchannel assignment and power control method that maximizes network energy efficiency while managing a large number of access requests. A distributed cooperative massive access approach based on DRL can meet the reliability and low latency requirements of URLLC in massive access scenarios [185]. Additionally, multi-access edge computing, as an emerging computing paradigm, has the capability to power large-scale IoT devices and novel mobile applications. How to deploy edge AI using limited computing resources in MEC environments is a research-worthy issue. Authors in [194] implement edge AI microservices by combining multiple dense models. This approach can further reduce deployment costs while meeting QoS constraints. For the massive deployment of network slices in 6G, researchers propose a system featuring distributed and AI-driven management and orchestration. This system can autonomously perform intelligent network management and orchestration, thereby enabling automated and scalable management of network slices [195].

### v4

@@ 3.4.3 Agentic AI-RAN
Agentic AI represents a significant, potentially revolutionary evolution for RAN management, shifting from traditional automation towards autonomous systems capable of pursuing complex goals with minimal human intervention [jiang2025largeaimodelsagentic, xiao2025agenticainetworking6g]. Operating under a principle of bidirectional interaction where AI empowers the network and the network enhances AI, agentic AI offers a paradigm for intelligent operation, administration, and maintenance in the complex environment of 6G networks. This approach enables systems to perceive, reason, decide, and act autonomously [dev2025advancedarchitecturesintegratedagentic]. Unlike conventional AI reliant on predefined rules, agentic AI RAN aims to optimize objectives like resource allocation, service assurance, and energy efficiency through dynamic adaptation to the network environment, thereby more effectively supporting emerging services such as autonomous intelligent IoT and embodied intelligence applications.
@@
The core capabilities underpinning agentic AI RAN include sophisticated multi source perception leveraging channel information, sensing data, and visual inputs to understand the network state [xiao2024llmagents6gorchestrator, xiao2025agenticainetworking6g]. Based on this perception, agents employ advanced reasoning processes, perhaps using chain of thought techniques to plan sequences of actions. They then interact with the network fabric by utilizing network function APIs as tools to execute the planned configurations. Crucially, effective decision making and planning are deeply reliant on integrated retrieval mechanisms. These provide agents dynamic access to essential knowledge bases containing 3GPP standards, historical logs, network policies, and other relevant information, enabling context aware reasoning and ensuring compliance [zhang2025agenticaigenerativeinformation].
@@
Building upon these capabilities, specific frameworks and architectural concepts are proposed to realize agentic AI-RAN. For example, AgentNet [xiao2025agenticainetworking6g] envisions specialized networking ecosystems supporting heterogeneous AI agents like foundation model agents and embodied agents, facilitating efficient information exchange and coordination. Generative foundation models themselves can serve as interactive knowledge resources or be specialized into agents capable of understanding tasks, orchestrating workflows, and managing network components through well defined interfaces [xiao2024llmagents6gorchestrator]. This agent driven architecture offers inherent simplifications, notably through control and user plane separation, and facilitates dynamic service orchestration and intent based networking, allowing high level goals to be translated into automated network adjustments [dev2025advancedarchitecturesintegratedagentic].
Similar to the hallucination issues in code generation, the stochastic nature of agentic AI introduces even more critical risks in operational environments. Specifically, agents may generate plausible but incorrect network configurations or decisions. For instance, an agent might hallucinate a resource allocation command that assigns users to a restricted frequency band (e.g., radar bands), violating spectrum regulations and causing severe interference to primary services. Besides, a temporary drop in traffic might be misinterpreted, leading to a critical macro base station being wrongly put into deep sleep. The result is an instant coverage gap and widespread call drops. To address these challenges, establishing a trustworthy agentic workflow is essential. First, leveraging RAG enables agents to cross-verify actions against up-to-date 3GPP standards and historical network logs, thereby reducing fact-conflicting hallucinations caused by outdated or missing domain knowledge [liu2024hallucinationawareoptimizationlargelanguage]. Second, integrating structured causal reasoning allows agents to explicitly construct causal directed acyclic logic graphs before execution. This ensures the agent captures the correct conditional dependencies between network variables, mitigating logically inconsistent hallucinations where generated actions violate physical propagation laws [li2025mitigatinghallucinationslargelanguage]. Finally, a separate AI verification layer serves as a critical safeguard. This validator model or rule-based engine audits all generated commands against physical network constraints, ensuring invalid configurations are intercepted before they can impact the live network [wang2025wirelesshallucinationgenerativeaienabled].
Beyond these reliability concerns, the performance overhead introduced by the API-based abstraction layer poses a significant challenge. Specifically, this overhead stems from the cumulative latency of model inference and the network round-trip time associated with API calling. This magnitude of delay makes direct agentic control unsuitable for ultra-low-latency data plane operations. To address this, agentic AI-RAN essentially requires a hierarchical control framework. In this architecture, the agent operates in the “slow loop” to translate intents and configure high-level policies, such as defining energy-saving strategies or network slicing parameters, while traditional tightly-coupled algorithms maintain the “fast loop” to guarantee millisecond-level responsiveness for critical tasks like real-time beamforming or packet scheduling. Furthermore, energy efficiency remains a critical constraint that warrants a profound analysis (as detailed in Section 6.3). Continued research is thus essential to develop computationally efficient AI models, secure agent interactions, and highly adaptive AI-driven processes tailored to the demanding operational realities of 6G [dev2025advancedarchitecturesintegratedagentic].
@@ 3.4.4 Multi-Agent Collaboration
The distributed and complex nature of 6G systems necessitates the use of multi agent systems where multiple autonomous AI agents interact and coordinate to achieve shared or individual goals [zhang2025multiagentreinforcementlearningwireless, jiang2025largeaimodelsagentic]. Agentic AI networking inherently involves collaboration as agents deployed across different network locations such as user equipment edge servers or base stations, therefore, they must work together for tasks like resource allocation interference management or service orchestration [xiao2025agenticainetworking6g]. Wireless distributed networks provide the necessary infrastructure platform for these multi agent interactions [zhang2025multiagentreinforcementlearningwireless].
@@
Multi-agent reinforcement learning (MARL) is a key enabling paradigm allowing agents to learn collaborative strategies through interaction. MARL frameworks address how agents learn optimal policies considering the actions and states of others moving from centralized training and execution towards more decentralized approaches like centralized training with decentralized execution or fully decentralized training and execution which better suit the distributed nature of 6G [zhang2025multiagentreinforcementlearningwireless]. Effective collaboration hinges on mechanisms for communication and information sharing considering network constraints. Techniques like graph enhanced MARL information bottleneck enhanced MARL and mirror learning are being explored to improve communication efficiency robustness and scalability in multi agent settings. Generative information retrieval can also support collaboration by providing agents access to shared knowledge bases or contextual information relevant to joint tasks [zhang2025agenticaigenerativeinformation].
@@
Structured workflows and specific agent roles are key to effective collaboration in practical agentic systems. For instance, LLM agents can automate physical layer tasks through a coordinated approach, with agents dedicated to task awareness, environmental observation, system configuration, and API invocation, all responding to high-level requests and environmental perception [xiao2024llmagents6gorchestrator]. Likewise, generative simulation frameworks employ agents for simulation generation, test design, execution, and result interpretation, collaborating via a central orchestrator to iteratively refine network models [rezazadeh2025generative6gsimulationexperimental]. Architectures like AgentNet further enhance inter-agent collaboration, supporting learning and knowledge transfer across diverse agent types [xiao2025agenticainetworking6g]. Crucially, agentic AI also facilitates human-AI collaboration, augmenting human capabilities in knowledge-intensive tasks and complex operations [jiang2025largeaimodelsagentic].
@@
While multi-agent collaboration holds immense promise, significant challenges must be addressed to fully realize its potential, particularly in areas like 6G networks. These challenges include managing communication overhead, ensuring effective coordination despite partial observability, handling non-stationarity arising from evolving agent policies, and maintaining scalability as agent numbers grow [zhang2025multiagentreinforcementlearningwireless]. Overcoming these hurdles and developing robust, efficient collaboration mechanisms is paramount to achieving the vision of truly intelligent and autonomous 6G networks.
3.5 Practical Framework for Applying Large AI Models
To bridge the gap between the extensive capabilities of large AI models and their practical application in wireless communications, a structured implementation framework is essential. This framework elucidates a systematic approach for practitioners to harness these models effectively.
3.5.1 Problem Formulation and Layer Identification
The process begins with problem formulation and layer identification, where the specific challenge, such as channel associated prediction or network resource management, is clearly defined and mapped to its corresponding communication layer [LLM4WM, NetLLM]. This initial step includes identifying the scope and objectives of the AI integration. For instance, tasks like transceiver design fundamentally reside at the physical layer [10104549], while network optimization and management are network layer concerns [10829820]. Semantic communications introduces its own layer of challenges and opportunities [gunduz2022beyond].
3.5.2 Model Selection and Adaptation
Following this, the framework proceeds to model selection and adaptation. This phase involves choosing an appropriate model architecture. For example, a practitioner might employ Transformers for their strength in handling sequential data in end to end communication systems [choukroun2022error], or leverage generative models like diffusion models for their capabilities in transceiver design [10104549]. For multimodal tasks such as ISAC, a MLLM might be the most suitable choice [cheng2025large]. The selected model then undergoes domain specific adaptation. Techniques such as parameter efficient fine tuning, including methods like LoRA, are crucial for tailoring the general knowledge of the model to the unique condition and constraints of wireless systems without necessitating complete retraining [hu2021lora, LLM4WM].
3.5.3 Data Integration and Multimodal Fusion
The third stage is data integration and multimodal fusion. Wireless systems generate a rich tapestry of data, including signals, images, and network logs. An effective framework must incorporate mechanisms to process and align these diverse data streams. This is exemplified in applications like ISAC, which fuses sensor data for enhanced environmental awareness [cheng2025large, li2025large]. It is also critical in network operations, where models analyze voluminous data sources such as base station logs and performance metrics to diagnose faults [10829820]. The ability to synthesize multimodal information enables more holistic and context aware decision making.
3.5.4 Interaction and Deployment
Finally, the framework culminates in interaction and deployment. This involves establishing how the AI model will interface with the network infrastructure. For many applications, this is realized through an agentic paradigm, where the large AI model acts as an intelligent agent that perceives the network state, reasons over complex data, and executes actions via application programming interfaces [xiao2025agenticainetworking6g, dev2025advancedarchitecturesintegratedagentic]. This approach, central to the concept of Agentic AI RAN, enables a dynamic and autonomous loop of monitoring, analysis, and control, paving the way for truly AI native wireless networks.
@@ 4 Wireless Communications for Large AI Models
Figure 12:   The outline of Section 4.
While Section 3 has demonstrated the transformative potential of WLAMs in optimizing wireless networks, realizing these applications in practice introduces a new set of challenges. The massive parameter scale that enables the superior performance of WLAMs simultaneously imposes immense pressure on computation, storage, and communication resources. Consequently, the primary research focus shifts from exploring the capacity of AI to optimize networks to addressing the network engineering challenges necessary to efficiently support large-scale AI models. To address these challenges for deployment and training within distributed wireless environments, this section explores the key wireless communication strategies and technologies that enable the efficient operation of large AI models, as illustrated in Fig. 12. To provide a clear and structured overview, our discussion is organized around four foundational pillars, each directly mapping enabling technologies to specific 6G performance indicators. Specifically, we examine how edge intelligence resolves computational and storage bottlenecks for distributed inference on constrained devices; how distributed collaborative learning (e.g., FL, SL) balances massive connectivity with stringent data privacy and training efficiency; how AirComp overcomes communication bottlenecks to minimize data aggregation latency and enhance spectral efficiency; and how physical layer security provides a lightweight, resilient defense against eavesdropping and jamming attacks. These pillars are then synthesized into a practical framework designed to guide the development of robust and efficient systems for supporting WLAM. The related works are summarized in Table 6.
@@
Edge Intelligence [lin2023pushing] Deploying LLMs at the 6G mobile edge for key applications like healthcare and robotics control in cloud-based deployment. Identify key challenges and discuss enabling techniques for efficient LLMs edge training and inference.
[zhang2025communication] Distributed on-device LLM inference using tensor parallelism. Propose a communication-efficient framework to accelerate the all-reduce operations in tensor parallelism.
[wdmoe] Splitting MoE LLMs with the gating network on an edge server and expert networks distributed on mobile devices. Propose the WDMoE architecture for collaborative inference, expert selection and bandwidth allocation to minimize latency.
DCML [OpenFedLLM] Collaborative and privacy-preserving training of LLMs on decentralized private data using FL. Propose OpenFedLLM, a concise, integrated, and research-friendly framework/codebase for FL on LLMs.
[shfMIOT] Federated fine-tuning and collaborative reasoning for large AI models at the network edge. Propose advanced FL paradigms like clustered, hierarchical, and asynchronous FL to address practical wireless challenges.
[zhao2024fedsllm] Efficiently fine-tuning LLMs over wireless networks by combining data and model parallelism. Propose the FedsLLM framework, which integrates LoRA with FSL to balance the workload between clients and server.
AirComp [liu2023over] Enabling distributed inference of large AI models that require non-linear aggregation functions like Max/Average pooling. Proposing a generalized AirComp framework to support non-linear functions in distributed large AI models.
[liu2024over] Efficiently fusing distributed sensor data for real-time inference in large multi-modal AI models. Propose an AirComp framework that exploits feature sparsity to enable scalable data aggregation for large perception models.
[shfTWC] Parameter-efficient federated fine-tuning of LLMs over wireless networks. Propose the AirFL-LORA framework using AirComp to achieve communication-efficient aggregation of low-rank gradient updates.
PLS [mitev2023physical] Providing a foundational security layer for distributed large AI model operations in 6G. Provide a comprehensive review of PLS for 6G, paving the way for intelligent networks with large AI models.
[infocomfederated] Defining the fundamental security limits for wireless federated learning, a key challenge for large AI models. Establish the theoretical limits for securing the high-dimensional parameter updates of large AI models.
[federatedPHY] Ensuring resilient federated training of LLM over wireless channels in the presence of physical layer attacks like jamming. Establishes a direct mathematical link between the training loss of a LLM and the communication MSE of the physical layer.
@@ 4.1 Edge Intelligence
6G networks are expected to support in-network, distributed AI capabilities at the edge, facilitating collaborative machine learning across devices with varying computational resources [lin2023pushing, 9797722, 10091939]. However, one of the main challenges is that many edge devices have limited computing power and storage capacity. To address this, wireless technologies are evolving to support the distributed large AI models [zeng2024implementation].
4.1.1 Distributed Collaborative Inference
While collaborative training is a key application, a critical challenge for deploying pre-trained LLMs is enabling collaborative inference across multiple resource-constrained devices. This avoids the need to store a massive model on a single node and leverages parallel processing to reduce latency. Recent research has focused on two primary paradigms for splitting a single large model for collaborative inference.
One prominent approach is based on architectural model splitting, particularly for MoE models. The authors in [wdmoe] propose the WDMoE framework, where an MoE-based LLM is split by placing the main network body and the gating network on an edge server, while distributing the individual expert networks across multiple mobile devices. This scheme effectively utilizes the parallel processing capabilities of edge devices for the expert layers, which constitute a significant portion of the model’s computation.
Another powerful paradigm is tensor parallelism, which partitions the internal weight matrices of an LLM across different devices. While this enables fine-grained parallel computation within each layer, it introduces a communication bottleneck due to the frequent ‘all-reduce’ operations required to aggregate intermediate results. To overcome this, the authors in [zhang2025communication] propose a communication-efficient framework that uses AirComp to accelerate this aggregation step, significantly improving the speed and scalability of on-device collaborative inference.
@@ 4.1.2 Resource Allocation in Heterogeneous Systems
Effective resource allocation in wireless networks is essential for supporting large AI model training and deployment across edge devices [10091939]. These collaborative inference paradigms transform the traditional resource allocation problem into a more complex, cross-layer optimization challenge. The key task evolves from merely allocating communication resources to the intelligent assignment of model components such as expert networks or tensor shards based on the heterogeneous capabilities of edge devices. This creates a new set of optimization problems, for instance, the joint expert selection and bandwidth allocation required in MoE-based systems to minimize latency [wdmoe], and the co-optimization of tensor shard placement and communication resources needed to prevent bottlenecks in tensor parallelism frameworks [zhang2025communication]. Successfully solving these emergent problems is central to realizing the full potential of collaborative edge intelligence for large AI models.